40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14516     Entry time: Fri Apr 5 00:33:58 2019     In reply to: 14503     Reply to this: 14519   14521
Author: gautam 
Type: Update 
Category: ALS 
Subject: Promising IR ALS noise 

Summary:

I set up a free-space beat on theNW side of the PSL table between the IR beam from the PSL and from EX, the latter brought to the PSL table via ~40m fiber. Initial measurements suggest very good performance, although further tests are required to be sure. Specifically, the noise below 10 Hz seems much improved.

Details:

Attachment #1 shows the optical setup. 

  • I used two identical Thorlabs F220APC collimators to couple the light back into free space, reasoning that the mode-matching would be easiest this way.
  • Only 1 spare K6Xs collimator mount was available (this has the locking nut on the rotational DoF), so I used a K6X for the other mount. The fast axis of the Panda fibers were aligned as best as possible to p-polarization on the table by using the fact that the connector key is aligned to the slow axis.
  • I cut the power coupled into the PSL fiber from ~2.6mW to ~880uW (using a HWP + PBS combo before the input coupling to the fiber) to match the power from EX.
  • The expected signal level from these powers and the NF1611 transimpedance of 700 V/A is ~320 mVpp. After alignment tweaking, I measured ~310mVpp (~ -5dBm) into a 50 ohm input on a scope, so the mode-matching which means the polarization matching and mode overlap between the PSL and EX beams are nearly optimal.
  • To pipe the signal to the delay line electronics, I decided to use the ZHL-3A (G=27dB, 1dB compression at 29.5dBm per spec), so the signal level at the DFD rack was expected (and confirmed via 50 ohm input on o'scope) to be ~19dBm.
  • This is a lot of signal - after the insertion loss of the power splitter, there would still be ~15dBm of signal going to the (nominally 10dBm) LO input of the demod board. This path has a Teledyne AP1053 at the input, which has 10dB gain and 1dBm compression at 26dBm per spec. To give a bit of headroom, I opted on the hacky solution of inserting an attenuator (5dB) in this path - a better solution needs to be implemented.
  • The differential outputs of the demod board go to the CDS system via an AA board (there is no analog whitening).

Yehonathan came by today so I had to re-align the arms and recover POX/POY locking. This alllowed me to lock the X arm length to the PSL frequency, and lock the EX green laser to the X arm length. GTRX was ~0.36, whereas I know it can be as high as 0.5, so there is definitely room to improve the EX frequency noise suppression.

Attachment #2 shows the ALS out-of-loop noise for the PSL+X combo. The main improvements compared to this time last year are electronic. 

  • The failed experiment of making custom I/F amplifier was abandoned and Rich Abbott's original design was reverted to. 
  • New power splitter was installed with 3dB less insertion loss.
  • According to the RF path level monitor, the signal level at the RF input to the demod board is ~10dBm. Per my earlier characterization, this will give us the pretty beefy frequency discriminant of ~15uV/Hz.
  • I estimate the frequency noise of the detection electronics + ADC noise now translate to 1/3 the frequency noise compared to the old system. With some analog whitening, this can be made even better, the electronics noise of the DFD electronics (~50nV/rtHz) is estimated to be <10mHz/rtHz equivalent frequency noise. 
  • Note that the calibration from phase-tracker-servo to units of Hz (~14 kHz / degree) was not changed in the digital system - this should only be a property of the delay line length, and hence, should not have changed as a result of the various electronics changes to the demod board and other electronics.

Next steps:

  • Improve pointing of green beam into X arm cavity.
  • I plan to recover the green beat note as well and digitize it using the second available DFD channel (eventually for the Y arm) - then we can simultaneously compare the the green and IR performance (though they will have different noise floors as there is less green light on the green beat PDs, and I think lower transimpedance too).
Quote:

Mix the beams in free space. We have the beam coming from EX to the PSL table, so once we mix the two beams, we can use either a fiber or free-space PD to read out the beatnote. 

  • This approach means we lose some of the advantages of the fiber based setup (e.g. frequent alignment of the free-space MM of the two interfering beams may be required).
  • Potentially increases sensitivity to jitter noise at the free-space/fiber coupling points
Attachment 1: IMG_7388.JPG  1.549 MB  Uploaded Fri Apr 5 01:58:08 2019  | Hide | Hide all
IMG_7388.JPG
Attachment 2: freeSpace_IR_beat.pdf  41 kB  Uploaded Fri Apr 5 01:58:15 2019  | Hide | Hide all
freeSpace_IR_beat.pdf
ELOG V3.1.3-